# Rapid, strongly magnetized accretion in the zero-net-vertical-flux shearing box

### v1

@@ 7.3 Application to global models
Some of the most important questions raised by this work relate to its application to global simulation results, e.g., those of Gaburov et al. (2012), H+24, and Guo et al. (2024). Gaburov et al. (2012) studied the disruption of a magnetized gas cloud by a black hole and found disks with midplane $\beta\sim 0.1$ supported almost entirely by magnetic pressure. The global multi-physics solutions of H+24, which form self-consistently from cosmological initial conditions via capture of gas from a cloud complex, evolve to an extremely strongly magnetized disk, with $\beta$ reaching ${\sim}10^{-4}$ in the midplane and large accretion rates $\alpha\gg 1$. Guo et al. (2024) studied the fueling of M87* and found that cooling of the hot intracluster gas produces a disk dominated by strong toroidal magnetic fields with $\beta\sim 10^{-2}$.
@@
Our results capture a number of general features of the Gaburov et al. (2012), H+24  and Guo et al. (2024) solutions, such as the dominant, single-polarity toroidal field and modestly sub-Alfvénic fluctuations. The vertical profile shape of the magnetic field and density found here are also very similar to those in H+24 and Guo et al. (2024). Most importantly, our results show that the magnetic flux does not immediately escape vertically, as previously argued from both local and global simulations (e.g., Salvesen et al., 2016a; Fragile & Sądowski, 2017). Rather, so long as the toroidal field is sufficiently strong ($\beta\lesssim 1$), the flux can sustain itself against escape via a dynamo mechanism.
@@
However, one key feature in which our solutions differ markedly from those of Gaburov et al. (2012), H+24  and Guo et al. (2024) is the midplane $\beta$, which is sustained at $\beta\ll 1$ in the global simulations but always collapses towards $\beta\sim 1$ given sufficient resolution in the local simulations presented here. This difference in midplane $\beta$ is particularly important in the Gaburov et al. (2012) and H+24 applications, since low $\beta$ leads to much lower gas densities, which helps the disk avoid fragmentation due to gravitational instability (a long-standing problem in fueling luminous AGN; Goodman 2003).
@@
Of course, since our local simulations also saturate at $\beta\ll 1$ when $H_{\rm th}$ is poorly resolved (the density collapses to the grid resolution; see Fig. 3), one possible cause of the difference is simply that the global cases are affected by resolution, artificially lowering $\beta$. However, it is also plausible that physics not included in the shearing box simulations presented here could be important and lead to different properties of the low-$\beta$ state in the simulations of Gaburov et al. (2012), H+24 and Guo et al. (2024). One is simply that the latter are all global: global instabilities and/or radial advection of field could be important for maintaining low $\beta$. In addition, H+24 in particular includes much more physics than we have here (e.g., multiphase thermodynamics, accretion from a dynamic interstellar medium leading to warps, radiation, and self gravity).
@@
Motivated by the limitations of the shearing box in capturing some potentially important physics, in a companion paper (Guo et al., in prep) we use Athenak simulations to study a simple model of the formation and evolution of a global disk initially dominated by a toroidal magnetic field; the disk forms from the inflow of a rotating gas cloud that has significant azimuthal magnetic flux. The global results in Guo et al. (in prep) are similar in many respects to the low-$\beta$ state presented here, although the midplane $\beta\sim 5$ is usually somewhat larger than that found in the shearing box and the midplane $\alpha\sim 0.1$ somewhat lower as a result (this difference may be at least partially resolution related; see App. B). As in our local simulations, the only circumstances in which Guo et al. (in prep) find a $\beta\ll 1$ midplane is when the thermal scale-height of the disk is not resolved. Although there are some interesting quantitative differences, this comparison of idealized local and global results supports the idea that the low-$\beta$ state is robust but that a midplane with $\beta\ll 1$ is not a generic outcome.
@@
The idealized global simulations in Guo et al. (in prep) and the local simulations presented here still differ dramatically in physics complexity from the simulations of H+24.101010This is less true for Gaburov et al. 2012 who studied a more idealized problem; indeed the idealized global model of Guo et al. (in prep) was chosen to be similar to Gaburov et al. (2012)’s calculation. It thus remains an open possibility that in some circumstances global simulations can realize a (resolved) $\beta\ll 1$ low-$\beta$ state, even though the local simulations and simplified global models do not. To speculate on the conditions under which at $\beta\ll 1$ may be sustained, we refer back to Fig. 13, which shows: (i) even as the system is losing net flux, it is regenerated near the midplane (just more slowly than it is produced) and (ii) that field is regenerated well above the midplane and sustains $\beta_{\rm mid}\ll 1$, so long as the midplane density collapse is halted (in this case, by the grid resolution). These observations hint that the cause of the loss of flux (leading to $\beta_{\rm mid}\sim 1$) could be the density collapse: in the absence of some particular scale height $H_{\rho}$ for the system to “choose,” it collapses to (nearly) $H_{\rm th}$. If so, then one might expect that any other physical process that halts the density collapse with $H_{\rho}\gg H_{\rm th}$ might also — as a side effect — halt the loss of flux. Another (non-exclusive) possibility is the opposite: that halting the loss of flux could stop the collapse of the density.
@@
As relates to the former possibility, there are a number of ways in which the density collapse could be halted in the global case, the most obvious of which is another source of supersonic midplane turbulence (it must be supersonic so that the turbulent pressure $P_{\rm turb}$ exceeds $P_{\rm th}$, thus leading to a larger scale-height and lower midplane density). As discussed extensively in H+24 and §4.1, there exist other global instabilities — not captured by the shearing box but in principle captured by the idealized global simulations of Guo et al. (in prep) — which can have faster growth rates than the standard MRI at $\beta\ll 1$ (Pessah & Psaltis, 2005; Das et al., 2018). These are driven by global currents and thus depend on the radial gradient of the (super-thermal) toroidal field, reducing to the local MRI when ${d\ln B_{\phi}}/{d\ln R}=-1$ such that $\bm{J}\propto\nabla\times(\overline{B_{\phi}}\hat{\bm{\phi}})\approx r^{-1}\partial_{r}(r\overline{B_{\phi}})\hat{\bm{z}}\approx 0$ (see §4.1 and Begelman & Armitage 2023). Such instabilities could plausibly produce supersonic turbulence in global simulations, and it is possible that they would be present in H+24 but not in Guo et al’s more idealized global simulations, given the different conditions in each. In support of this idea, we note that the $B_{\phi}$ gradient in H+24’s simulation is close to, but not exactly, $\hat{B}=-1$ (see their figure 3), as would be expected if the instability was driving the system close to marginality (see also their figure 16). If correct, this implies the existence of yet another switch-like transition between accretion states, with the system choosing between (i) ultra-low $\beta$ (active global instabilities with Parker dynamo maintaining $\beta\ll 1$), (ii) low $\beta$ (as studied here in the shearing box, with $\beta_{\rm mid}\sim 1$), and (iii) high-$\beta$ cycles. The choice between these would be determined primarily by the total azimuthal and/or radial magnetic flux available, as well as ${d\ln B_{\phi}}/{dR}$.
@@
In addition to the above scenario, there also exist a variety of other potential physical explanations for the maintenance of $\beta\ll 1$ in a global context. Perhaps the most natural is the possibility that an additional source of $\overline{\bm{B}}$, sufficient to overwhelm its vertical loss (see Fig. 13), would stop the density’s collapse by maintaining $\beta\ll 1$ (i.e., the reverse causation to that discussed above). This could be a simple consequence of radial flux advection, as discussed extensively in H+24, which would add additional growth and loss terms into the $\partial_{t}\overline{B_{y}}$ equation (see Fig. 12), depending on the profile and relative compression of $B_{\phi}$ and $B_{R}$ as they are transported in the disk. Indeed, because of the large accretion rates, the accretion timescales in the H+24 simulation can be extremely short, of order several orbital times, and thus potentially shorter than the timescale over which the flux is lost vertically in our local simulations. Such radial advection seems insufficient to maintain $\beta\ll 1$ at the midplane in Guo et al. (in prep) but these simulations explored just one realization of a disk dominated by strong toroidal fields; other global contexts may advect flux inwards more efficiently.
@@
Overall, clearly more work on this subject is needed. While global setups will undeniably be the most relevant for study of these issues, aspects of the Pessah & Psaltis (2005) instabilities may be possible to capture in local domains. The importance of thermodynamics, more complex initial conditions, and other physics present in H+24 and Guo et al. (2024) should also be explored to see if they influence the magnetization in the low-$\beta$ state. Whatever the outcome, there clearly exists a rich array of behaviors in strongly magnetized disks, in particular various “switches” — state transitions that depend on the magnetic field strength and geometry, which flip the system between weakly and strongly accreting states. This has interesting applications for our understanding of quasars, X-ray binaries, CVs, and other high-energy accreting systems.
@@ Acknowledgments
We thank Omer Blaes, Charles Gammie, Minghao Guo, Matt Kunz, Eve Ostriker, Jim Stone, Yashvardhan Tomar, and especially Yuri Levin for useful conversations. JS acknowledges the support of the Royal Society Te Apārangi, through Marsden-Fund grant MFP-UOO2221 and Rutherford Discovery Fellowship RDF-U001804. This work was also supported in part by Simons Investigator grants from the Simons Foundation (PH, EQ). This research was part of the Frontera computing project at the Texas Advanced Computing Center, which is made possible by National Science Foundation award OAC-1818253. Further computational support was provided by the New Zealand eScience Infrastructure (NeSI) high performance computing facilities, funded jointly by NeSI’s collaborator institutions and through the NZ MBIE. This research benefited from interactions at the Kavli Institute for Theoretical Physics, supported by NSF PHY-2309135.
@@ Appendix A Dependence on box size
Figure 16: Testing the effect of changing the horizontal box dimensions on the transition to the low-$\beta$ state, as diagnosed by the time evolution of $B_{y,{\rm mid}}/B_{y0}$ in the main panel (see Fig. 4; plots of $\beta_{\rm mid}$ or $\alpha$ show similar features). The system appears robust to the choice of $L_{x}$, as indicated by $\beta$R0.1-narrow reaching the low-$\beta$ state, as well as to longer box lengths in the azimuthal direction ($\beta$R0.1-long). But, decreasing $L_{y}$ by a factor of 2 in $\beta$R0.1-short and $\beta$R0.1-short-narr (both with $L_{y}=L_{z})$ causes an otherwise identical system to lose its flux and transition to the high-$\beta$ state. At lower $\beta$ and smaller $H_{\rm th}$ it can maintain the low-$\beta$ state even at $L_{y}=L_{z}$ ($\beta$R0.05-short). The inset shows the azimuthal wavenumber $k_{\gamma,{\rm max}}$ of the fastest growing Parker mode (blue; Eq. 9) and MRI mode (red; $k_{y}\approx\Omega/v_{{\rm A}y}$) as a function of $z$ for the profile of the $\beta$R0.1 run, normalized by $k_{y0}=2\pi/L_{y}$. The dotted lines show the first three wavenumbers that fit in the box.
@@
Given our hypothesis that MRI is necessary for maintaining turbulence in the midplane, while the Parker instability drives stronger turbulence and regenerates the field above, our tentative conclusion is that scales several times larger than that of the fastest-growing Parker quasi-interchange mode are needed to sustain the low-$\beta$ state. This is qualitatively consistent with the arc-merging picture seen in 2D Parker-instability simulations in Johansen & Levin (2008), as well as the fact that the Parker scale (9) depends on $H_{\rm th}$ while that of the MRI does not (explaining the results of $\beta$0.05-short). However, we cannot rule out other possible explanations at the present time and a more detailed study is needed.
@@ Appendix B Dependence on resolution
We explore the dependence of the results on numerical resolution via a set of simulations at successively decreasing resolution with the same parameters as $\beta$0.1-mr (see table 2). This reveals interesting non-monotonic behavior, illustrated in the top panel of Fig. 17, whereby the transport level decreases with resolution up to some point (around $N_{z}=56$ here), then suddenly increases again to $\alpha\gtrsim 1$. This behavior can be explained by appealing to the phenomenology of the $\beta\ll 1$ state described in §5.6. If the resolution is sufficiently high that the thermal density profile can be resolved, $\alpha$ decreases with resolution because the midplane turbulence amplitude decreases with resolution [^11] Note that this is the reverse of the behavior seen at high resolution in the high-$\beta$ state, where the turbulence amplitude drops with increasing resolution (Ryan et al., 2017); but, given that the low-$\beta$ state involves larger-scale fluctuations and has a stronger mean field (see Fig. 2), this is not unexpected. (not shown; the turbulence amplitude in the corona is not a strong function of resolution). As the turbulence level drops, the density in the midplane collapses to the thermal profile across a wider range in $z$, and thus starts to resemble the high-$\beta$ state, reaching $\beta>1$ in the midplane in the lower-resolution cases. This is illustrated in the lower panel of Fig. 17 for $\beta$0.1-84 (cf. Fig. 3). However, at resolutions below ${\simeq}1.5$ zones per $H_{\rm th}$ for this setup ($\beta$0.1-56 has 1.7 zones per $H_{\rm th}$), the density collapse is halted by the grid itself, maintaining a $\beta\ll 1$ midplane as a side effect. The midplane turbulence also maintains a larger amplitude, perhaps via overshoot from the Parker-unstable atmosphere, thus causing much larger transport (the $\beta$0.1-32 simulation also has stronger outflows and has lost around half its density by $t\Omega/2\pi=40$).
@@
This behavior is broadly consistent with that discussed in the main text, although there do also appear to be important effects related to $L_{z}/H_{\rm th}$ and/or the initial conditions. In simulation $\beta$0.001, which has ${\simeq}1.1$ zones zones per $H_{\rm th}$ like $\beta$0.1-32 but a much smaller $H_{\rm th}/L_{z}$, the density collapse is also supported by the resolution and it maintains a high-transport $\beta\ll 1$ midplane. However, $\beta$0.01 and $\beta$0.01-hr, which have resolutions per $H_{\rm th}$ similar to $\beta$0.1-112 and $\beta$0.1-224 respectively, maintain $\beta<1$ in the midplane, albeit with a nearly thermal density profile across a wider range of $z/H_{\rm th}$ and a flatter $P_{\rm B}$ profile (see Figs. 3 and 10c). Similarly $\beta$0.1-tall has a transport level somewhat less than half that of $\beta$0.1-224, which has the same resolution per $H_{\rm th}$ but a smaller vertical extent (the different boundary conditions may also be important in this difference, since we see from table 1 that Outflow conditions
@@
Figure 18: Panels a-b show the effect changing the density floor below the default value of $\rho_{\rm flr}=10^{-4}$ by a factor of ${\simeq}3$ ($\beta$0.1-lr-lowflr; $\rho_{\rm flr}=3\times 10^{-5}$) and $10$ ($\beta$0.1-lr-lowerflr; $\rho_{\rm flr}=10^{-5}$). Panel (a) shows $\alpha(t)$, while (b) shows the time-averaged density and pressure profiles, as in Fig. 17. The horizontal dotted lines illustrate the density floor in each case. Panels (c)-(d) show the similar test of restarting $\beta$0.1-mr at $t\Omega=400$ with a lower floor $\rho_{\rm flr}=3\times 10^{-5}$.

### v3

@@ 7.3 Application to global models and $\beta\ll 1$ fields
Some of the most important questions raised by this work relate to its application to global simulation results, e.g., those of Gaburov et al. (2012), H+24, and Guo et al. (2024). Gaburov et al. (2012) studied the disruption of a magnetized gas cloud by a black hole and found disks with midplane $\beta\sim 0.1$ supported almost entirely by magnetic pressure. The global multi-physics solutions of H+24, which form self-consistently from cosmological initial conditions via capture of gas from a cloud complex, evolve to an extremely strongly magnetized disk, with $\beta$ reaching ${\sim}10^{-4}$ in the midplane and large accretion rates $\alpha\gg 1$. Guo et al. (2024) studied the fueling of M87* and found that cooling of the hot intracluster gas produces a disk dominated by strong toroidal magnetic fields with $\beta\sim 10^{-2}$. Other global simulations have likely also yielded similar states, for instance Machida et al. (2006), who observed a $\beta\sim 0.1$ magnetically supported solution emerge from a cooling disk that contracted vertically, or Sądowski (2016), who reported a thermally stable strongly magnetized disk arising from idealized initial conditions with a net radial field.
@@
Our results capture a number of general features of these solutions, such as the modestly sub-Alfvénic fluctuations and dominant, single-polarity toroidal field. While local simulations with NVF also create strong toroidal fields and some outwardly similar features, the different morphology observed in global simulations with NVF (e.g., Zhu & Stone, 2018; Mishra et al., 2020), suggests that the results of H+24, Guo et al. (2024), and Gaburov et al. (2012) do not result from “patches” of vertical field driving the MRI (moreover, these simulations do not appear to involve a vertical field that is at all coherent; see, e.g., H+24 figure 6). The vertical profile shape of the magnetic field and density found here are also very similar to those in H+24 and Guo et al. (2024).
@@
However, one key feature in which our solutions differ markedly is the midplane $\beta$: while some of these global disks sustain $\beta\ll 1$ in the midplane, our local simulations always collapse to $\beta\sim 1$ given sufficient resolution. This difference in midplane $\beta$ is particularly important in understanding AGN, since low $\beta$ leads to much lower gas densities, which helps the disk avoid fragmentation due to gravitational instability (a long-standing problem in fueling luminous AGN; Goodman 2003).
@@
Given that our local simulations also saturate at $\beta\ll 1$ when $H_{\rm th}$ is poorly resolved (the density collapses to the grid resolution; see Fig. 4), one possible cause of the difference is simply that the global cases are affected by resolution, artificially lowering $\beta$. However, it is also plausible that physical processes not captured in the shearing box simulations presented here could play a significant role, leading to differences. Clearly, our simulations omit global effects, and we discuss below how mechanisms such as global instabilities or radial advection of flux may be needed to sustain low $\beta$. Similarly, H+24 in particular incorporates a much wider set of physical processes — e.g., multiphase thermodynamics, accretion from a dynamic interstellar medium (which induces warps), radiation, and self gravity — which may also have relevance.
@@
Motivated by the limitations of the shearing box in capturing some potentially important physics, in a companion paper (Guo et al., in prep) we use Athenak simulations to study a simple model of the formation and evolution of a global disk initially dominated by a toroidal magnetic field, which forms from the inflow of a rotating gas cloud seeded with azimuthal magnetic flux. The global results in Guo et al. (in prep) are similar in many respects to the low-$\beta$ state presented here and seem to exhibit a similar resolution dependence, showing a $\beta\ll 1$ midplane only when the thermal scale-height of the disk is not resolved (see App. B). This comparison of idealized local and global results supports the idea that the low-$\beta$ state is robust but that a midplane with $\beta\ll 1$ is not a generic outcome.
@@
The idealized global simulations in Guo et al. (in prep) and the local simulations presented here still differ dramatically in physics complexity from the simulations of H+24.121212This is less true for Gaburov et al. 2012 who studied a more idealized problem; indeed the idealized global model of Guo et al. (in prep) was chosen to be similar to Gaburov et al. (2012)’s calculation. It thus remains an open possibility that in some circumstances global simulations can realize a (resolved) $\beta\ll 1$ low-$\beta$ state, even though the local simulations and simplified global models do not. To speculate on the conditions under which $\beta\ll 1$ may be sustained, we refer back to Fig. 13, which shows: (i) even as the system is losing net flux, it is regenerated near the midplane (just lost more slowly than it is produced) and (ii) that field is regenerated well above the midplane and sustains $\beta_{\rm mid}\ll 1$, so long as the midplane density collapse is halted (in this case, by the grid resolution). These observations hint that the cause of the loss of flux (leading to $\beta_{\rm mid}\sim 1$) could be the density collapse: in the absence of some particular scale height $H_{\rho}$ for the system to “choose,” it collapses to (nearly) $H_{\rm th}$. If so, then one might expect that any other physical process that halts the density collapse with $H_{\rho}\gg H_{\rm th}$ might also — as a side effect — halt the loss of flux. Another (non-exclusive) possibility is the opposite: that halting the loss of flux could stop the collapse of the density.
@@
As relates to the former possibility, there are a number of ways in which the density collapse could be halted in the global case. Other than radiation, which can provide another non-thermal pressure that supports the midplane against collapse (e.g., Jiang & Blaes, 2020), the most obvious possibility is supersonic turbulence in the midplane (it must be supersonic so that the turbulent pressure $P_{\rm turb}$ exceeds $P_{\rm th}$, thus leading to a larger scale-height and lower midplane density). As discussed extensively in H+24 and §4.1, there exist other global instabilities that can have faster growth rates than the standard MRI at $\beta\ll 1$ (Pessah & Psaltis, 2005; Das et al., 2018). These are driven by global currents and thus depend on the radial gradient of the (super-thermal) toroidal field, reducing to the local MRI when ${d\ln B_{\phi}}/{d\ln R}=-1$ such that $\bm{J}\propto\nabla\times(\overline{B_{\phi}}\hat{\bm{\phi}})\approx r^{-1}\partial_{r}(r\overline{B_{\phi}})\hat{\bm{z}}\approx 0$ (see §4.1 and Begelman & Armitage 2023). Such instabilities could plausibly produce supersonic turbulence in global simulations, and it is possible that they would be present in H+24 but not in Guo et al’s more idealized global simulations, given the different conditions in each. In support of this idea, we note that the $B_{\phi}$ gradient in H+24’s simulation is close to, but not exactly, $\hat{B}=-1$ (see their figure 3), as would be expected if the instability was driving the system close to marginality (see also their figure 16). If correct, this implies the existence of yet another switch-like transition between accretion states, with the system choosing between (i) ultra-low $\beta$, with active global instabilities and the Parker dynamo maintaining $\beta\ll 1$, (ii) low $\beta$, as studied here in the shearing box, with $\beta_{\rm mid}\sim 1$, and (iii) high-$\beta$ cycles. The choice between these would be determined primarily by the total azimuthal and/or radial magnetic flux available, as well as ${d\ln B_{\phi}}/{dR}$.
@@
In addition to the above scenario, the maintenance of $\beta\ll 1$ in a global context could be explained by various other effects, from radiation pressure and multi-phase gas structure to an additional source of magnetic flux. The latter could stop the density’s collapse by maintaining $\beta\ll 1$ (i.e., the reverse causation to that discussed above) if it were sufficient to overwhelm vertical field losses. Such a source could arise as a simple consequence of radial flux advection, as discussed extensively in H+24, adding additional growth and loss terms into the $\partial_{t}\overline{B_{y}}$ equation (10), which depend on the profile and relative compression of $B_{\phi}$ and $B_{R}$ as they are transported in the disk. Indeed, because of the large accretion rates, the accretion timescales in the H+24 simulation can be extremely short, of order several orbital times, and thus potentially shorter than the timescale over which the flux is lost vertically in our local simulations. While such radial advection seems insufficient to maintain $\beta\ll 1$ at the midplane in Guo et al. (in prep), these simulations explored just one class of idealized setup; other global solutions may advect flux inwards more efficiently.
@@
Overall, clearly more work on this subject is needed. While global setups will undeniably be the most relevant for study of these issues, aspects of the Pessah & Psaltis (2005) instabilities may be possible to capture in local domains. The importance of thermodynamics, more complex initial conditions, and other physics present in various global studies should also be explored to see if they influence the magnetization in the low-$\beta$ state. Whatever the outcome, there clearly exists a rich array of behaviors in strongly magnetized disks, in particular various magnetic “switches” — state transitions that depend on the magnetic-field strength and geometry. This has interesting applications for our understanding of quasars, X-ray binaries, CVs, and other high-energy accreting systems.
@@ Acknowledgments
We thank Omer Blaes, Charles Gammie, Minghao Guo, Matt Kunz, Eve Ostriker, Jim Stone, Yashvardhan Tomar, and especially Yuri Levin for enlightening discussion, as well as the anonymous referee, whose useful suggestions improved the manuscript. JS acknowledges the support of the Royal Society Te Apārangi, through Marsden-Fund grant MFP-UOO2221 and Rutherford Discovery Fellowship RDF-U001804. This work was also supported in part by Simons Investigator grants from the Simons Foundation (PH, EQ) and by NSF AST grant 2107872. This research was part of the Frontera computing project at the Texas Advanced Computing Center, which is made possible by National Science Foundation award OAC-1818253. Further computational support was provided by the New Zealand eScience Infrastructure (NeSI) high performance computing facilities, funded jointly by NeSI’s collaborator institutions and through the NZ MBIE. This research benefited from interactions at the Kavli Institute for Theoretical Physics, supported by NSF PHY-2309135.
@@ Appendix A Dependence on box size
Figure 16: Testing the effect of changing the horizontal box dimensions on the transition to the low-$\beta$ state, as diagnosed by the time evolution of $B_{y,{\rm mid}}/B_{y0}$ in the main panel (see Fig. 5; plots of $\beta_{\rm mid}$ or $\alpha$ show similar features). The system appears robust to the choice of $L_{x}$, as indicated by $\beta$R0.1-narrow reaching the low-$\beta$ state, as well as to longer box lengths in the azimuthal direction ($\beta$R0.1-long). But, decreasing $L_{y}$ by a factor of 2 in $\beta$R0.1-short and $\beta$R0.1-short-narr (both with $L_{y}=L_{z})$ causes an otherwise identical system to lose its flux and transition to the high-$\beta$ state. At lower $\beta$ and smaller $H_{\rm th}$ it can maintain the low-$\beta$ state even at $L_{y}=L_{z}$ ($\beta$R0.05-short). The inset shows the azimuthal wavenumber $k_{\gamma,{\rm max}}$ of the fastest growing Parker mode (blue; Eq. 9) and MRI mode (red; $k_{y}\approx\Omega/v_{{\rm A}y}$) as a function of $z$ for the profile of the $\beta$R0.1 run, normalized by $k_{y0}=2\pi/L_{y}$. The dotted lines show the first four wavenumbers that fit in the box.
@@
Given our hypothesis that MRI is necessary for maintaining turbulence in the midplane, while the Parker instability drives stronger turbulence and regenerates the field above, our tentative conclusion is that scales several times larger than that of the fastest-growing Parker quasi-interchange mode are needed to sustain the low-$\beta$ state. This is qualitatively consistent with the arc-merging picture seen in 2D Parker-instability simulations in Johansen & Levin (2008) as well as the fact that the Parker scale (9) depends on $H_{\rm th}$ while that of the MRI does not (thus, the MRI alone would not explain the results of $\beta$0.05-short). However, we cannot rule out other possible explanations at the present time and a more detailed study is needed.
@@ Appendix B Dependence on resolution
We explore the dependence of the results on numerical resolution via a set of simulations with successively fewer grid points and the same parameters as $\beta$0.1-mr (see table 2). This reveals interesting non-monotonic behavior, illustrated in the top panel of Fig. 17; the transport level decreases with coarsening resolution up to some point (around $N_{z}=56$ here), then suddenly increases again to $\alpha\gtrsim 1$. This behavior can be explained by appealing to the phenomenology of the $\beta\ll 1$ state described in §5.5. If the resolution is sufficiently fine that the thermal density profile can be resolved, $\alpha$ decreases with coarsening resolution because the midplane turbulence amplitude decreases [^13] Note that this is the reverse of the behavior seen at high resolution in the high-$\beta$ state, where the turbulence amplitude drops with finer resolution (Ryan et al., 2017). Presumably this difference arises because the low-$\beta$ state involves larger-scale fluctuations and has a stronger mean field (see Fig. 3). (not shown; the turbulence amplitude in the corona is not a strong function of resolution). As the turbulence level drops, the density in the midplane collapses to the thermal profile across a wider range in $z$, and thus starts to resemble the high-$\beta$ state, reaching $\beta>1$ in the midplane in poorly resolved cases. This is illustrated in the lower panel of Fig. 17 for $\beta$0.1-84 (cf. Fig. 4). However, at resolutions below ${\simeq}1.5$ zones per $H_{\rm th}$ for this setup ($\beta$0.1-56 has 1.7 zones per $H_{\rm th}$), the density collapse is halted by the grid itself, maintaining a $\beta\ll 1$ midplane as a side effect. The midplane turbulence also maintains a larger amplitude, perhaps via overshoot from the Parker-unstable atmosphere, thus causing much larger transport (the $\beta$0.1-32 simulation also has stronger outflows and has lost around half its density by $t\Omega/2\pi=40$).
@@
This behavior is broadly consistent with that discussed in the main text, although there do also appear to be important effects related to $L_{z}/H_{\rm th}$ and/or the initial conditions. In simulation $\beta$0.001, which has ${\simeq}1.1$ zones per $H_{\rm th}$ like $\beta$0.1-32, but with a much smaller $H_{\rm th}/L_{z}$, the density collapse is also supported by the resolution and it maintains a high-transport $\beta\ll 1$ midplane. However, $\beta$0.01 and $\beta$0.01-hr, which have resolutions per $H_{\rm th}$ similar to $\beta$0.1-112 and $\beta$0.1-224 respectively, maintain $\beta<1$ in the midplane, albeit with a nearly thermal density profile across a wider range of $z/H_{\rm th}$ and a flatter $P_{\rm B}$ profile (see Figs. 4 and 10c). Similarly $\beta$0.1-tall has a transport level somewhat less than half that of $\beta$0.1-224, which has the same resolution per $H_{\rm th}$ but a smaller vertical extent (the different boundary conditions may also be important in this difference, since we see from table 1 that Outflow conditions
@@
Figure 18: Panels a-b show the effect of changing the density floor below the default value of $\rho_{\rm flr}=10^{-4}$ by a factor of ${\simeq}3$ ($\beta$0.1-lr-lowflr; $\rho_{\rm flr}=3\times 10^{-5}$) and $10$ ($\beta$0.1-lr-lowerflr; $\rho_{\rm flr}=10^{-5}$). Panel (a) shows $\alpha(t)$, while (b) shows the time-averaged density and pressure profiles, as in Fig. 17. The horizontal dotted lines illustrate the density floor in each case. Panels (c)-(d) show the similar test of restarting $\beta$0.1-mr at $t\Omega=400$ with a lower floor $\rho_{\rm flr}=3\times 10^{-5}$.